OpenStudy (anonymous):

the value of √-9 . √-16

OpenStudy (anonymous):

\[\sqrt{-9}-\sqrt{-16}=\sqrt{(-1)(9)}-\sqrt{(-1)(16)}=\sqrt{-1}\sqrt{9}-\sqrt{-1}\sqrt{16}=??\] Can you finish it off? @JNaruto5796

OpenStudy (anonymous):

?????? @JNaruto5796

OpenStudy (anonymous):

Greatest Common Factor (GCF) Factoring by Grouping Sum of Cubes Difference of Perfect Squares Difference of Cubes Factoring Trinomials Grouping Help!!! please